OpenStudy (anonymous):

A formula to find the volume of a cone is V=1/3πr^2h where: V=Volume r=radius h=height which of the following shows the volume of a cone formula solved for h? A)h=3Vπr^2 B)h=1/3Vπr^2 C)h=3V/πr^2 h=3/V πr^2

OpenStudy (anonymous):

\[V=1/3\pi r^2h\]\[3V=\pi r^2h\]\[3V/\pi r^2=h\] C